# Causal Inference with Missing Exposures and Missing Outcomes

### v2

@@ Discussion
Here, we outline an approach to account for the dependence of participant outcomes within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [67, 68, 69, 70]). By following the Causal Roadmap or a similar framework for causal inference [21, 71], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [51, 70, 25, 26, 72]).
@@
Suppose it is reasonable to assume that participant outcomes are dependent within households, but effectively independent between households. Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. (See, for example, [70].) Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [51, 26, 72].
@@
If clustering is present, estimation and inference must be adjusted. First, the cross-validation scheme used within Super Learner must respect the independent unit. Concretely, participants in a given cluster are all assigned to the same sample-split. Second, for influence curve-based inference, let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$ [18]. Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience. Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$. We then proceed with variance estimation using the cluster-level influence curve. This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference.

### v3

@@ Discussion
Appendix S4: The Delta Method for Ratio-Estimands
The Delta method is commonly applied to derive influence curve-based inference for asymptotically linear estimators.[79] Recall our exposure-specific estimand is given by the following ratio:
Following the notation of Moore and van der Laan,[45] we denote the numerator with $\mu_{1}$ and the denominator with $\mu_{0}$. Let $IC_{1}$ and $IC_{0}$ be the corresponding influence curves. Then based on the Delta method, the influence curve for $log[\Psi(\mathbb{P};a)]$ is given by $IC(a)=1/\mu_{1}\times IC_{1}-1/\mu_{0}\times IC_{0}$. We obtain a variance estimate with the sample variance of $IC(a)$ scaled by sample size $N$. Using that variance estimate, we conduct hypothesis testing and create 95% confidence intervals before exponentiating to transform back to the original scale.
Now consider the relative association:
$RR=\frac{\Psi(\mathbb{P};1)}{\Psi(\mathbb{P};0)}=\frac{\mu_{1}/\mu_{0}}{\mu_{3}/\mu_{2}}$ (4)
with corresponding subscripts for the influence curves of the TMLEs. Then the influence curve for the $log(RR)$ is given by
$IC_{RR}=(-1/\mu_{0})\times IC_{0}+(1/\mu_{1})\times IC_{1}+(1/\mu_{2})\times IC_{2}-(1/\mu_{3})\times IC_{3}$ (5)
We obtain inference as above.
Appendix S5: Simulation Study
We conducted a simulation study designed to mimic the causal model given in Figure 3 in the main text. The data generating process is as follows where $\text{expit}(x)=(1+e^{-x})^{-1}$.
  • Baseline covariates: $L_{0,1}\sim\text{Bernoulli}(0.5)$, $L_{0,2}\sim\text{Uniform}(0,1)$, and $L_{0,3}\sim\text{Uniform}(0,1)$.
  • Measurement indicator for the exposure: $\Delta_{A}\sim\text{Bernoulli}(\text{expit}(1.7+0.4L_{0,1}-0.5L_{0,3}))$
  • Exposure: $A\sim\text{Bernoulli}(\text{expit}(-0.3+1.5L_{0,1}-1.2L_{0,2}))$, observed only when $\Delta_{A}=1$.
  • Measurement indicator for the baseline outcome: $\Delta_{Y_{0}}\sim\text{Bernoulli}(\text{expit}(2.0+0.3L_{0,1}-0.4L_{0,2}-0.3A+0.2\Delta_{A}))$
  • Baseline outcome: $Y_{0}\sim\text{Bernoulli}(\text{expit}(-2.6+1.2L_{0,1}+0.8L_{0,2}-0.5L_{0,3}+0.7A+0.3A\cdot L_{0,1}))$, observed only when $\Delta_{Y_{0}}=1$.
  • Time-varying covariates: $L_{1,1}=\text{clip}(\mathcal{N}(0.3A+0.2L_{0,1}+0.15L_{0,2},\ 0.25^{2}),0,1)$ and $L_{1,2}\sim\text{Bernoulli}(\text{expit}(-0.3+1.2A+0.8L_{0,1}-0.6L_{0,3}+0.5Y_{0}))$
  • Measurement indicator for the endline outcome: $\Delta_{Y_{1}}\sim\text{Bernoulli}(\text{expit}(1.1+5.5L_{1,1}^{2}A-5.0L_{0,3}L_{1,1}L_{1,2}-4.0Y_{0}A+0.5L_{0,1}))$, among those with $\Delta_{Y_{0}}=1$ and equal to 0 otherwise.
  • Endline outcome: $Y_{1}\sim\text{Bernoulli}(\text{expit}(-2.0+5.0A\cdot L_{1,1}^{2}-4.5L_{0,3}^{2}L_{1,2}+1.5L_{0,1}+1.2A))$, observed only when $\Delta_{Y_{1}}=1$.
We calculated the true value of the exposure-specific estimands $\Psi(\mathbb{P};1)$ and $\Psi(\mathbb{P};0)$ as well as the relative and absolute effects by generating $10^{6}$ counterfactuals and taking the relevant means, ratios, and differences, respectively.
We implemented TMLE, G-computation, and inverse probability weighting (IPW) for the exposure-specific estimands and the effects. Mirroring the real-data analysis, we used Super Learner within TMLE to combine predictions from generalized linear models, the mean, and multivariate adaptive regression splines with and without covariate screening. We used parametric regressions in the singly robust approaches. Also mirroring the real-data analysis, we implemented a “Naïve-adjusted” approach: subset on participants with complete data ($\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$) and at-risk at baseline ($Y_{0}=0$) and then implemented TMLE with adjustment for confounding. For completeness in the simulations, we also implemented a “Naïve-unadjusted” approach: subset on participants with complete data ($\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$) and at-risk at baseline ($Y_{0}=0$) and then take the empirical mean outcome within exposure groups. Influence curve-based inference was used for variance estimation and confidence interval construction. We ran 1000 simulation studies at a sample size of $N=20000$ to reflect the motivating example.
Appendix S4.1: Simulation Results
The true values of the estimands were as follows: the expected outcome under the exposure $\Psi(\mathbb{P};1)=0.510$, the expected outcome under no exposure $\Psi(\mathbb{P};0)=0.177$, the risk ratio $RR=2.879$, and the risk difference $RD=0.333$. Table 1 provides results in terms of bias (average deviation between the point estimate and truth), empirical standard deviation (standard deviation [SD] of the point estimate across simulation replicates), the ratio of bias to empirical SD, and 95% confidence interval coverage (proportion of calculated intervals containing the truth). We report the bias-to-SD ratio as a diagnostic for the asymptotic behavior of our estimators: since the standard deviation shrinks at rate $\sqrt{N}$ while systematic bias remains fixed, a non-negligible bias-to-SD ratio indicates performance will deteriorate as sample size grows.
For all estimands, TMLE with Super Learner achieved nominal coverage, had negligible bias, and had bias-to-SD ratios well below 0.2. In contrast, G-computation and IPW only achieved nominal coverage for the expected outcome under the exposure $\Psi(\mathbb{P};1)$, but not the other estimands. Indeed, the coverage of G-computation was 93.2% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 91.6% for the risk ratio, and 92.1% for the risk difference. The coverage of IPW was only 12.3% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 19.7% for the risk ratio, and 58.9% for the risk difference.
Table 1: Over 1000 simulation iterations, estimator performance for TMLE, G-computation (“G-comp”), inverse probability weighting (IPW), and the naïve approach with and without confounding adjustment.
Estimand Estimator Bias Emp SD Bias/SD Coverage
$\Psi(\mathbb{P};1)$ TMLE 0.0014 0.0091 0.1560 0.9440
$\Psi(\mathbb{P};1)$ G-comp 0.0053 0.0083 0.6435 0.9500
$\Psi(\mathbb{P};1)$ IPW -0.0029 0.0089 -0.3231 0.9690
$\Psi(\mathbb{P};1)$ Naïve-adjusted 0.0565 0.0088 6.4347 0.0000
$\Psi(\mathbb{P};1)$ Naïve-unadjusted 0.1089 0.0080 13.5559 0.0000
$\Psi(\mathbb{P};0)$ TMLE 0.0003 0.0058 0.0578 0.9530
$\Psi(\mathbb{P};0)$ G-comp -0.0024 0.0054 -0.4546 0.9320
$\Psi(\mathbb{P};0)$ IPW 0.0179 0.0063 2.8376 0.1230
$\Psi(\mathbb{P};0)$ Naïve-adjusted 0.0102 0.0060 1.7023 0.6290
$\Psi(\mathbb{P};0)$ Naïve-unadjusted -0.0151 0.0052 -2.9191 0.0380
Risk Ratio TMLE 0.0056 0.1068 0.0525 0.9510
Risk Ratio G-comp 0.0733 0.1012 0.7247 0.9160
Risk Ratio IPW -0.2765 0.0953 -2.9025 0.1970
Risk Ratio Naïve-adjusted 0.1484 0.1054 1.4071 0.7240
Risk Ratio Naïve-unadjusted 0.9454 0.1316 7.1833 0.0000
Risk Difference TMLE 0.0011 0.0107 0.1012 0.9460
Risk Difference G-comp 0.0078 0.0097 0.7978 0.9210
Risk Difference IPW -0.0208 0.0109 -1.9127 0.5890
Risk Difference Naïve-adjusted 0.0463 0.0104 4.4549 0.0100
Risk Difference Naïve-unadjusted 0.1240 0.0095 13.0900 0.0000
We emphasize that TMLE, G-computation, and IPW rely on the same identification result. Therefore, differences in their performance are due to differences in statistical estimation. To explore further, Table 2 provides their performance for the numerator and denominator of our ratio-estimand:
$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$
TMLE had negligible bias for all components; its bias-to-SD ratio remained below 0.2. In contrast, G-computation exhibited meaningful bias, as indicated by the bias-to-SD ratio, for all components except for the numerator under $A=1$. The numerator/denominator decomposition identifies the source IPW’s poor performance: the numerator for $A=0$ is substantially biased with a bias-to-SD ratio of $2.79$. This bias propagates directly into meaningful bias and poor coverage for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, the risk ratio, and the risk difference.
Table 2: Further details on estimator performance with TMLE, G-comp, and IPW for the exposure-specific components of the ratio-estimands.
Estimator Exposure Component Truth Mean Est Bias Emp SD Bias/SD
TMLE $A=1$ Numerator 0.3593 0.3602 0.0009 0.0071 0.1274
TMLE $A=1$ Denominator 0.7047 0.7045 -0.0002 0.0056 -0.0346
TMLE $A=0$ Numerator 0.1499 0.1501 0.0002 0.0050 0.0419
TMLE $A=0$ Denominator 0.8464 0.8460 -0.0004 0.0044 -0.0957
G-comp $A=1$ Numerator 0.3593 0.3604 0.0011 0.0066 0.1686
G-comp $A=1$ Denominator 0.7047 0.6996 -0.0051 0.0054 -0.9529
G-comp $A=0$ Numerator 0.1499 0.1473 -0.0026 0.0046 -0.5549
G-comp $A=0$ Denominator 0.8464 0.8435 -0.0029 0.0044 -0.6663
IPW $A=1$ Numerator 0.3593 0.3571 -0.0021 0.0069 -0.3099
IPW $A=1$ Denominator 0.7047 0.7045 -0.0002 0.0056 -0.0344
IPW $A=0$ Numerator 0.1499 0.1649 0.0151 0.0054 2.7942
IPW $A=0$ Denominator 0.8464 0.8460 -0.0004 0.0045 -0.0950
Now we consider the naïve approaches, which are targeting a different (incorrect) estimand. Therefore, there was a substantial “causal gap”, even when adjusting for confounding.[53, 15, 14, 21, 48] By simply subsetting on participants with complete data and baseline risk, the naïve approaches failed to appropriately account for differential missingness. As shown in Table 1, the naïve approaches exhibited meaningful bias for all estimands. The coverage of naïve approach with confounding adjustment was 0% for the expected outcome under the exposure $\Psi(\mathbb{P};1)$, 62.9% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 72.4% for the risk ratio, and 1% for the risk difference. Finally, the coverage of naïve approach without confounding adjustment was <<5% for all estimands.
Altogether, our simulation results support the robustness of TMLE with Super Learner in settings resembling our applied analysis. Our simulation results further suggest the practical differences between TMLE, G-computation, and IPW in SEARCH-TB are due to TMLE’s flexible adjustment for confounding and differential missingness. Again, all three approaches are targeting the same estimand. Therefore, the simulation and applied results highlight how theoretical properties (e.g., double robustness) translate into practical impact.
@@
Here, we outline an approach to account for the dependence of participant outcomes within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [23, 24, 32, 6]). By following the Causal Roadmap or a similar framework for causal inference,[53, 27] we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [78, 6, 4, 50, 49]).
@@
Suppose it is reasonable to assume that participant outcomes are dependent within households, but effectively independent between households. Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. (See, for example, [6].) Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit.[78, 50, 49]
@@
If clustering is present, estimation and inference must be adjusted. First, the cross-validation scheme used within Super Learner must respect the independent unit. Concretely, participants in a given cluster are all assigned to the same sample-split. Second, variance estimation must account for clustering. Following Schnitzer et al.,[68] let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$. Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience.[68] Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$. We then proceed with variance estimation using the cluster-level influence curve. This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference. (As detailed [9], a slightly different formulation is needed if the causal estimand weights clusters, versus individuals, equally.)
Appendix S7: Missingness DAGs
There is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”).[43, 44] In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link.[46] They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures.[46, 85] Holovchak et al. extended this work for the effects of longitudinal exposures, while questioning the practical utility of the m-DAG approach.[30] Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis”.[30] As an alternative, we demonstrate how the Causal Roadmap, an established framework for causal inference, can used to define, identify, and estimate causal estimands, including the average treatment effect under missing exposures and Counterfactual Strata Effects.